# Bayesian Quantile Regression with Subset Selection: A Posterior Summarization Perspective

### v2

@@ Abstract
Quantile regression is a powerful tool for inferring how covariates affect specific percentiles of the response distribution. Existing methods either estimate conditional quantiles separately for each quantile of interest or estimate the entire conditional distribution using semi- or non-parametric models. The former often produce inadequate models for real data and do not share information across quantiles, while the latter are characterized by complex and constrained models that can be difficult to interpret and computationally inefficient. Further, neither approach is well-suited for quantile-specific subset selection. Instead, we pose the fundamental problems of linear quantile estimation, uncertainty quantification, and subset selection from a Bayesian decision analysis perspective. For any Bayesian regression model, we derive optimal and interpretable linear estimates and uncertainty quantification for each model-based conditional quantile. Our approach introduces a quantile-focused squared error loss, which enables efficient, closed-form computing and maintains a close relationship with Wasserstein-based density estimation. In an extensive simulation study, our methods demonstrate substantial gains in quantile estimation accuracy, variable selection, and inference over frequentist and Bayesian competitors. We apply these tools to identify the quantile-specific impacts of social and environmental stressors on educational outcomes for a large cohort of children in North Carolina.
62G08,
62F15,
62P25 ,
Variable selection,
decision theory,
interpretable machine learning,
Bayesian inference,
keywords:
[class=MSC]
keywords:
volume: TBAissue: TBA
@@
[…]
 [^4] An R package, QRSubsets, implementing the proposed methods is available online: https://github.com/jfeldman396/QRSubsets
1 Introduction
Quantile regression estimates the functional relationship between covariates and specific percentiles of a response variable. We focus on linear quantile regression, which assumes that the $\tau$th conditional quantile for a random variable $Y$ changes linearly as a function of $p$-dimensional predictors $\boldsymbol{x}$:
$Q_{\tau}\{Y\mid\boldsymbol{x},\boldsymbol{\beta}(\tau)\}=\boldsymbol{x}^{\intercal}\boldsymbol{\beta}(\tau).$ (1)
Estimated across $\tau$, the coefficients $\boldsymbol{\beta}(\tau)$ summarize how the covariates affect not only the location, but also the shape of the response distribution $Y\mid\boldsymbol{x}$. This is useful when interest lies in providing a more robust and comprehensive view of the relationship between covariates and the response variable. Consequently, quantile regression methods have been been applied in diverse settings, including medicine (Kottas and Gelfand, 2001), finance (Bassett and Chen, 2002), and environmental studies (Pandey and Nguyen, 1999), among many others.
The unique insight from linear quantile regression comes from identifying predictors with heterogeneous effects $\beta_{j}(\tau)\neq\beta_{j}(\tau^{\prime})$ for some $\tau\neq\tau^{\prime}$. This capability is essential when the covariates affect higher order moments or tails of $Y\mid\boldsymbol{x}$. For instance, in childhood educational outcome studies (Section 5), it is essential to determine whether socioeconomic variables, environmental exposures, and other key factors differentially impact low, medium, or high-achieving students, which can provide far-reaching implications for policy interventions.
Broadly, there are several important components and considerations in quantile regression. First, quantile-specific linear coefficient estimates are obtained to detect potentially heterogeneous covariate effects, including both magnitude and direction. Second, quantile-specific uncertainty quantification provides important context for these coefficients and facilitates comparisons across both quantiles and variables. Third, when $p$ is moderate or large, quantile-specific subset selection provides more interpretable and parsimonious summaries and identifies the most impactful covariates across the distribution $Y\mid\boldsymbol{x}$. Each of these targets must simultaneously respect the fundamental smoothness across quantiles: estimates, uncertainties, and selections should be similar for adjacent quantiles. Finally, the algorithms that deliver these results should be scalable in both the number of observations $n$ and the number of covariates $p$.
We first review existing Bayesian and frequentist methods for quantile regression, both to showcase the successes in this area and to highlight the need for methodological advances. We then introduce our approach and explain how it addresses each of the challenges above.
1.1 Separate Quantile Regressions
Bayesian inference proceeds by placing priors on $\boldsymbol{\beta}(\tau)$ and inferring posterior distributions separately for each $\tau$. Under a flat prior on $\boldsymbol{\beta}(\tau)$ and the likelihood (3), the maximum a posteriori estimator yields the same solution as (2). Broader theoretical justification is provided by Sriram et al. (2013), who detail sufficient conditions under which the posterior for $\boldsymbol{\beta}(\tau)$ is strongly consistent. Convenient parameter expansions have been developed to facilitate posterior sampling under (3) (Kozumi and Kobayashi, 2011; Fasiolo et al., 2021), which has led to widespread implementation with open source software (Benoit and Van den Poel, 2017).
In general, Bayesian quantile regression with the AL likelihood faces significant limitations. First, there is no information-sharing among nearby quantiles, which results in excessively large posterior uncertainties and underpowered inference for $\boldsymbol{\beta}(\tau)$, especially for extreme quantiles near zero or one (see Sections 4-5). Second, an AL likelihood must be specified for each $\tau$, which induces distinct Bayesian models for the same data, typically with no attempt to reconcile or combine them. Finally, the AL likelihood often produces a substantially inadequate model for real data across many, if not all $\tau$, which undermines the interpretability of the resulting inferences. Kowal and Wu (2023) proposed a transformation-based generalization of (3) to improve model adequacy, but did not address the previous two concerns.
1.2 Simultaneous Quantile Regression Methods
1.3 Subset Selection in Quantile Regression
Among separate quantile regression methods, modifications to (2) have been developed for quantile-specific variable selection. Sparse estimates are obtained by appending a penalty term to (2), such as an $\ell_{1}$-penalty (Wu and Liu, 2009; Belloni and Chernozhukov, 2011; Wang et al., 2012; Lee et al., 2014). The sparsity among penalized quantile regression estimates is controlled by a tuning parameter, typically selected via cross-validation, which can be computationally intensive for the (penalized) objective (2). These methods do not provide uncertainty quantification, and are limited in that common sparsity penalties i) introduce overshrinkage of nonzero effects and ii) severely restrict the subset search path. By design, these methods focus on selecting a single ”best” subset. However, even for moderate $p$ with correlated covariates or weak signals, there are often many subsets that offer similar predictive accuracy. Thus, it can be misleading to report only a single ”best” subset; consideration of a broader collection of ”near-optimal” subsets can be more comprehensive and informative, but requires a robust subset search (Kowal, 2022).
For Bayesian variable selection with separate AL likelihoods (3), it is common to use sparsity or shrinkage priors for $\boldsymbol{\beta}(\tau)$ (Li and Zhu, 2008; Alhamzawi et al., 2012; Chen et al., 2013). However, these priors do not resolve the fundamental inadequacies of the AL model. Further, each of these methods selects variables marginally, either via posterior inclusion probabilities or credible intervals that exclude zero. Under the AL likelihood, marginal posteriors for $\ \beta_{j}(\tau)$ are often characterized by large uncertainty, which results in severely underpowered variable selection (see Section 4).
1.4 Quantile Regression Through Posterior Summarization
  1. Fit a Bayesian regression model $\mathcal{M}$
  2a. Quantile estimation and uncertainty quantification
  2b. Subset search, filtration, and selection
Finally, we leverage the decision analysis framework to achieve quantile-specific subset search and selection. Our approach yields closed-form linear coefficient estimates for any quantile and any subset, and unlocks decision analysis strategies for variable selection previously deployed only for mean regression (Hahn and Carvalho, 2015; Woody et al., 2021; Kowal, 2021, 2022). We extend these approaches for quantile regression and design a subset search procedure to accumulate an acceptable family of subsets that provide strong predictive accuracy for each quantile. From this acceptable family, we recommend one subset based on the parsimony principle (i.e., the smallest subset in this family), but also construct variable importance metrics to avoid the overreliance on any single ”best” subset. This is especially important when multiple subsets offer similar predictive accuracy. Our variable importance metrics seek to provide additional context in this common scenario.
The remainder of the paper is organized as follows. In Section 2, we introduce posterior decision analysis for quantile regression. Section 3 describes our quantile-specific subset search and selection procedure. We provide a simulation study in Section 4 to evaluate the proposed methodology for prediction, inference, selection, and quantile crossing. In Section 5 we conduct a quantile regression analysis using data on end-of-grade test scores for children in North Carolina. We conclude in Section 6.
2 Bayesian Decision Analysis for Quantile Regression
Thus, (6) provides linear quantile estimation, specific to each quantile $\tau$ and each subset $S$, under the model $\mathcal{M}$ and the loss $\mathcal{L}$. We note that the proposed decision analysis framework is not strictly limited to linear actions $\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)$: it is possible to extract quantile-specific trees or additive functions of $\boldsymbol{x}$, akin to Woody et al. (2021) for mean regression. Here, we focus on linear actions in accordance with (1).
2.1 Point Estimation for Linear Quantiles
Lemma 2.1.
with zeros for indices $j\notin S$, where $\hat{Q}_{\tau}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}$, $\boldsymbol{\hat{Q}}_{\tau}(\boldsymbol{X})=\{\hat{Q}_{\tau}(\boldsymbol{x}_{1}),\dots\\ ,\hat{Q}_{\tau}(\boldsymbol{x}_{n})\}^{\intercal}$, and ${\boldsymbol{X}}_{S}$ the $n\times\lvert S\rvert$ matrix of active covariates for subset $S$.
Proof.
It suffices to observe that $E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\lVert Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})-\boldsymbol{x}_{i}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\rVert_{2}^{2}=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\lVert\{Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})-\hat{Q}_{\tau}({\boldsymbol{x}}_{i})\}+\{\hat{Q}_{\tau}({\boldsymbol{x}}_{i})-\boldsymbol{x}_{i}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\}\rVert_{2}^{2}=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\lVert Q_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})-\hat{Q}_{\tau}({\boldsymbol{x}}_{i})\rVert_{2}^{2}+\\ E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\lVert\hat{Q}_{\tau}({\boldsymbol{x}}_{i})-\boldsymbol{x}_{i}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\rvert_{2}^{2}$ where the first term is finite and does not depend on $\boldsymbol{\delta}_{S}(\tau)$. The remaining steps constitute an ordinary least squares solution. ∎
Corollary 1.
Proof.
With $\hat{s}(\boldsymbol{x}_{i})=E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\sigma$ for any $\boldsymbol{x}_{i}$, it suffices to observe that under (4), ${\hat{Q}}_{\tau}({\boldsymbol{x}_{i}})=\hat{\beta}_{1}(\tau)+\sum_{j=2}^{p}{x_{ij}}\hat{\beta}_{j}$. In addition, for the action using the full set of covariates, the covariate matrix is ${\boldsymbol{X}}$. Plugging this into (8) yields the result. ∎
Corollary 2.
Proof.
The result is immediate using the argument in Corollary 1 and observing that $\boldsymbol{\hat{Q}}_{\tau}({\boldsymbol{X}})={\boldsymbol{X}}\hat{\boldsymbol{\beta}}(\tau)$. ∎
While Corollaries 1 and 2 confirm reasonable behavior for the optimal action (6) under linear regression models, we emphasize the utility of Lemma 2.1 for quantile regression estimation under any Bayesian regression model. In particular, the analyst can prioritize curation of $\mathcal{M}$ to capture complex features of the entire conditional distribution $Y\mid\boldsymbol{x}$, including nonlinearity, skewness, heteroscedasticity, while (8) provides the optimal linear approximation of the model-based quantiles under the loss function (7).
2.2 Posterior Uncertainty Quantification
Corollary 3.
Proof.
For the homoscedastic linear regression, $Q_{\tau}(\boldsymbol{Y}\mid\boldsymbol{X},\boldsymbol{\theta})=[\{\beta_{1}+\sigma F^{-1}(\tau)\}+\sum_{j=2}^{p}x_{ij}\beta_{j}]_{i=1}^{n}$. Therefore, for the full set of covariates and any quantile, (9) is equal in distribution to the regression coefficients under the model $\mathcal{M}$ posterior, with only the intercept term varying as a function of the quantile. ∎
2.3 Connection with the Wasserstein Geometry
[…]
Lemma 2.2.
Proof.
It suffices to observe that objective in (13) may be expanded as $\sum_{\tau=\tau_{1}}^{\tau_{\ell}}E_{\boldsymbol{\theta}\mid\boldsymbol{y}}\sum_{i=1}^{n}\lVert{Q}_{\tau}(Y_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\theta})-\boldsymbol{x}_{i}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\rVert_{2}^{2}$, and since the optimization is unconstrained for $\boldsymbol{\delta}_{S}(\tau)$ across $\tau$, the summands may be optimized separately. Applying Lemma 2.1 separately for each $\tau_{m}$ yields the result. ∎
We emphasize that the relaxation of the density requirement maintains alignment with our primary goals: estimation, uncertainty quantification, and selection of quantile-specific regression coefficients. We do not claim to provide valid density estimates. We do, however, observe that the estimated quantiles tend to preserve probabilistic coherence, such as quantile non-crossing (see the supplementary material (Feldman and Kowal, 2024)). Furthermore, this relaxation does not affect the validity of the proposed decision analysis: we specify a carefully-chosen loss function and minimize the posterior expected loss under the model $\mathcal{M}$. Indeed, quantile regression often focuses on a few select quantiles $\tau$, such as upper or lower extremes, quartiles, and medians. Consideration of a fine grid $\tau\in\{\tau_{1},\dots,\tau_{\ell}\}$, while feasible with the proposed approach, may offer little additional information.
3 Quantile-Specific Subset Search and Selection
3.1 Subset Search with Decision Analysis
which resembles (14) but instead inherits a posterior distribution via $p(\boldsymbol{\theta}\mid\boldsymbol{y})$ under model $\mathcal{M}$. Thus, (15) enables uncertainty quantification for the performance of each subset. The evaluation criteria in (14) and (15) are used in conjunction to search and subsequently filter to the ”near-optimal” or acceptable family of subsets (Kowal, 2021), specific to each quantile $\tau$.
The inputs for the BBA algorithm are i) the model-based fitted quantiles $\{\hat{Q}_{\tau}(\boldsymbol{x}_{i})\}$, ii) the covariates $\{\boldsymbol{x}_{i}\}$, and iii) the maximum number of subsets $m_{k}$ to return for each subset size $k$, which should be chosen as large as computationally feasible. In our simulation and application, we set $m_{k}=50$. An efficient implementation of the BBA algorithm is available in the leaps package in R, which provides fast filtering for $p\leq 35$; see the supplement (Feldman and Kowal, 2024) for details on our model-assisted pre-screening approach when $p>35$.
3.2 Subset Filtration
3.3 Quantifying Variable Importance
A principal advantage of curating the quantile-specific acceptable family (17) is that it contains more information than any single (”best”) subset. In the common applied setting with moderate $p$, correlated covariates, or weak signals, there are typically many subsets that offer similar accuracy. Thus, we de-emphasize the selection of a single ”best” subset and instead provide more comprehensive summaries of the quantile-specific acceptable family.
4 Simulation Study
4.1 Simulation Design
By design, (20) yields linear conditional quantile functions $Q^{*}_{\tau}(\boldsymbol{x})=\boldsymbol{x}^{\intercal}\boldsymbol{\beta}^{*}(\tau)$, where $\beta^{*}_{1}(\tau)=2+\Phi^{-1}(\tau)$ and $\beta^{*}_{j}(\tau)=\xi^{*}_{j}+\Phi^{-1}(\tau)\gamma^{*}_{j},j=2,\ldots,p$, with $\Phi$ the standard normal CDF. The covariate values $\boldsymbol{x}_{i}$ are simulated independently for $i=1,\ldots,n$ from a Gaussian copula with uniform marginals and copula correlation $\rho_{\ell j}=0.5^{\lvert\ell-j\rvert}$, which provides varying degrees of correlation among the predictors. Additional simulations with independent covariates ($\rho_{\ell j}=0$) are available in the supplement.
We carry out our simulation design for $(n,p)\in\{(500,20),(200,50)\}$ and $\mbox{HetRatio}\in\{0.5,1\}$ across 50 independent repetitions of the data generating process in each setting.
with model parameters $\boldsymbol{\theta}=(\sigma,\boldsymbol{\xi},\boldsymbol{\gamma})$. By design, both the mean and variance functions are misspecified under $\mathcal{M}$: both include all covariates, even those with true null effects, while the variance function is nonlinear. The full Bayesian specification of $\mathcal{M}$ is available in the supplement, but we note here that regularizing priors are specified on both $\boldsymbol{\xi}$ and $\boldsymbol{\gamma}$ and the model is estimated using STAN (Carpenter et al., 2017). Posterior samples of the $\tau$th conditional quantile function under $\mathcal{M}$ are easily computed under (21): for each posterior sample $\boldsymbol{\theta}^{s}$ and any covariate value $\boldsymbol{x}$, we simply compute $Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta}^{s})=\boldsymbol{x}^{\intercal}\boldsymbol{\xi}^{s}+\sigma^{s}\exp(\boldsymbol{x}^{\intercal}\boldsymbol{\gamma}^{s})\Phi^{-1}(\tau)$.
Under each $(n,p,\mbox{HetRatio})$ combination, we curate acceptable families for $\tau\in\{0.01,0.05,0.25,0.5,0.75,0.95,0.99\}$ using the proposed methodology in Sections 2-3 under default settings ($m_{k}=50,\varepsilon=0.05$). Among the subsets in each $\mathbbm{A}_{0.05}(\tau)$, we evaluate quantile prediction, inference, and variable selection for $\mbox{S}_{small}(\tau)$.
[…]
4.2 Quantile Prediction
We first summarise the predictive performance of the competing methods on hold-out datasets $(\boldsymbol{X}_{test},\boldsymbol{y}_{test})$ with $n_{test}=1000$, which are generated independently and identically distributed to the training data. We compute out-of-sample quantile predictions, say $\tilde{Q}_{\tau}(\boldsymbol{X}_{test})$, for each competitor. For each $\tau$, we evaluate the performance using three metrics: i) the mean-squared error (MSE) between the ground-truth $Q_{\tau}^{*}(\boldsymbol{X}_{test})$ and predictions $\tilde{Q}_{\tau}(\boldsymbol{X}_{test})$, ii) the average check loss between between $\boldsymbol{y}_{test}$ and $\tilde{Q}_{\tau}(\boldsymbol{X}_{test})$, and iii) the calibration of $\tilde{Q}(\boldsymbol{X}_{test})$ computed via $n_{test}^{-1}\sum_{i=1}^{n_{test}}\mathbbm{1}\{y_{test_{i}}\leq\tilde{Q}_{\tau}(\boldsymbol{x}_{test_{i}})\}$, which should be near $\tau$ for a well-calibrated quantile estimator (see the supplementary material (Feldman and Kowal, 2024)). The results for MSE and check loss for $\tau\in\{0.01,0.5,0.99\}$ and $n=200,p=50,\mbox{HetRatio}=1$ across simulations are presented in Figure 1; results for other simulation settings are qualitatively similar (Feldman and Kowal, 2024).
4.3 Uncertainty Quantification
The intervals are evaluated for calibration, measured by the empirical coverage, and sharpness, measured by the average interval widths, in Table LABEL:covrates. Narrow intervals that achieve the 95% nominal coverage are preferred. Most notably, the posterior actions for both $\mbox{S}_{full}$ and $\mbox{S}_{small}(\tau)$ yield significantly more narrow intervals than $\mbox{AL}_{Bayes}$, often by a factor of 2-5. These effects are especially pronounced for extreme quantiles near zero or one, where $\mbox{AL}_{Bayes}$ is excessively conservative and underpowered. Importantly, both $\mbox{S}_{full}$ and $\mbox{S}_{small}(\tau)$ maintain the nominal coverage in nearly all settings. As expected, the sparsity of $\mbox{S}_{small}(\tau)$ produces the most narrow intervals, but also sacrifices empirical coverage for any active variables excluded from $\mbox{S}_{small}(\tau)$. The results for the other simulation settings are similar, and can be found in the supplementary material (Feldman and Kowal, 2024).
Table 3: $\boldsymbol{n=200,p=50,\textbf{\mbox{HetRatio}}=1}$
$\tau$ 0.01 0.05 0.25 0.5 0.75 0.95 0.99
Coverage Rate $\mbox{S}_{small}(\tau)$ 0.83 0.86 0.88 0.90 0.88 0.85 0.83
$\mbox{S}_{full}$ 0.98 0.97 0.95 0.94 0.95 0.97 0.97
$\mbox{AL}_{Bayes}$ 0.98 0.97 0.94 0.94 0.95 0.97 0.98
$\mbox{S}_{full}$ 3.56 2.87 2.02 1.76 2.03 2.90 3.60
$\mbox{AL}_{Bayes}$ 8.64 4.95 3.10 2.73 3.04 4.99 9.19
Coverage rates and average widths of 95% credible intervals for the posterior action (9) with $\mbox{S}_{small}(\tau)$ or $\mbox{S}_{full}$ as well as $\mbox{AL}_{Bayes}$. The intervals for $\mbox{S}_{small}(\tau)$ and $\mbox{S}_{full}$ are significantly more narrow than those for $\mbox{AL}_{Bayes}$, especially for quantiles near zero or one, and typically maintain nominal coverage. $\mbox{S}_{small}(\tau)$ sacrifices some coverage in favor of sparsity, and thus provides the most narrow intervals.
4.4 Selection
We evaluate quantile-specific subset selection by computing true positive rates (TPR) and true negative rates (TNR) for the variables selected by $\mbox{S}_{small}(\tau)$, aLASSO and $\mbox{AL}_{Bayes}$. Here, $\mathcal{M}$ is not a competitor: it does not specify linear quantiles or any mechanism for quantile-specific variable selection. The results are presented in Table LABEL:Table1.
Across all settings, $\mbox{S}_{small}(\tau)$ provides a superior balance between TPR and TNR. $\mbox{AL}_{Bayes}$ is underpowered and overconservative due to the excessively wide posterior credible intervals that are used for selection (see Table LABEL:covrates). The frequentist competitor aLASSO often produces similar TPRs but lower TNRs compared to $\mbox{S}_{small}(\tau)$, and thus selects too many variables.
[…]
Table 4: $\boldsymbol{n=200,p=50,\textbf{\mbox{HetRatio}}=1}$
$\tau$ 0.01 0.05 0.25 0.5 0.75 0.95 0.99
TPR $\mbox{S}_{small}(\tau)$ 0.43 0.54 0.69 0.79 0.63 0.57 0.46
aLASSO 0.21 0.76 0.89 0.95 0.89 0.71 0.27
$\mbox{AL}_{Bayes}$ 0.00 0.02 0.22 0.35 0.23 0.03 0.00
TNR $\mbox{S}_{small}(\tau)$ 0.88 0.85 0.82 0.84 0.83 0.85 0.87
aLASSO 0.88 0.55 0.29 0.27 0.28 0.54 0.83
$\mbox{AL}_{Bayes}$ 0.97 0.97 0.97 0.96 0.97 0.97 0.93
Table 5: $\boldsymbol{n=500,p=20,\textbf{\mbox{HetRatio}}=1}$
$\tau$ 0.01 0.05 0.25 0.5 0.75 0.95 0.99
TPR $\mbox{S}_{small}(\tau)$ 0.96 0.98 0.95 1.00 0.92 0.99 0.96
aLASSO 0.65 0.99 0.99 1.00 0.98 0.99 0.72
$\mbox{AL}_{Bayes}$ 0.00 0.31 0.83 0.99 0.86 0.36 0.00
TNR $\mbox{S}_{small}(\tau)$ 0.94 0.93 0.93 0.91 0.92 0.93 0.94
aLASSO 0.87 0.71 0.58 0.52 0.55 0.72 0.82
$\mbox{AL}_{Bayes}$ 1.00 1.00 0.99 0.99 0.99 1.00 1.00
True positive rates (TPR) and true negative rates (TNR) for variable selection averaged across simulations. The proposed approach ($\mbox{S}_{small}(\tau)$) neatly balances TPR and TNR. The Bayesian competitor ($\mbox{AL}_{Bayes}$) is significantly underpowered due to excessively wide posterior credible intervals, while the frequentist competitor (aLASSO) often overselects variables (low TNRs).
Childhood educational outcomes are affected by adverse environmental exposures such as air quality and lead, and social stressors like poverty and racial residential isolation (Miranda et al., 2007). Using a large cohort of children in North Carolina (CEHI, 2020), we analyze student-level 4th end-of-grade (EoG) reading test scores ($y_{i}$) along with individual birth information, mother/child demographics, air quality exposure, blood lead measurements, and socioeconomic factors ($\boldsymbol{x}_{i}$). The variables are summarised in Table 4. Previous analyses of these data have exclusively used homoscedastic mean regression for estimation and variable selection (Kowal et al., 2021; Kowal, 2022; Bravo et al., 2022). However, it is essential to know which, if any, of the environmental exposures, social stressors, or other factors are heterogeneous; i.e., to determine whether (any of) these effects vary in their impact on low, medium, or high-achieving students. Thus, quantile regression is an informative tool—especially with the ability to quantify effect directions and magnitudes, measure and report uncertainty, and provide quantile-specific subset selection.
@@ 5 Quantile Estimation, Inference, and Selection for Educational Outcomes
RI Residential isolation, at time of EoG test

### v3

@@ Abstract
Quantile regression is a powerful tool in epidemiological studies where interest lies in inferring how different exposures affect specific percentiles of the distribution of a health or life outcome. Existing methods either estimate conditional quantiles separately for each quantile of interest or estimate the entire conditional distribution using semi- or non-parametric models. The former often produce inadequate models for real data and do not share information across quantiles, while the latter are characterized by complex and constrained models that can be difficult to interpret and computationally inefficient. Further, neither approach is well-suited for quantile-specific subset selection. Instead, we pose the fundamental problems of linear quantile estimation, uncertainty quantification, and subset selection from a Bayesian decision analysis perspective. For any Bayesian regression model, we derive optimal and interpretable linear estimates and uncertainty quantification for each model-based conditional quantile. Our approach introduces a quantile-focused squared error loss, which enables efficient, closed-form computing and maintains a close relationship with Wasserstein-based density estimation. In an extensive simulation study, our methods demonstrate substantial gains in quantile estimation accuracy, variable selection, and inference over frequentist and Bayesian competitors. We use these tools to identify and quantify the heterogeneous impacts of multiple social stressors and environmental exposures on educational outcomes across the full spectrum of low-, medium-, and high-achieving students in North Carolina.
Variable selection, decision theory, interpretable machine learning, Bayesian inference
1 Introduction
Quantile regression estimates the functional relationship between covariates and specific percentiles of a response variable. In linear quantile regression, the $\tau$th conditional quantile for a random variable $Y$ is modeled as a function of $p$-dimensional predictors $\boldsymbol{x}$ via
$Q_{\tau}\{Y\mid\boldsymbol{x},\boldsymbol{\beta}(\tau)\}=\boldsymbol{x}^{\intercal}\boldsymbol{\beta}(\tau)$ (1)
(see Section 2.4 for the nonlinear case). Estimated across $\tau$, the coefficients $\boldsymbol{\beta}(\tau)$ summarize how the covariates affect not only the location, but also the shape of the response distribution $Y\mid\boldsymbol{x}$. The unique insight from linear quantile regression comes from identifying predictors with heterogeneous effects $\beta_{j}(\tau)\neq\beta_{j}(\tau^{\prime})$ for some $\tau\neq\tau^{\prime}$. For instance, when analyzing education data (see Sections 1.1 and 5), it is important not only to identify the factors that impact educational outcomes, but also to determine whether those effects are different for low-, medium, or high-achieving students. Such heterogeneous, quantile-specific effects can have far-reaching implications for policy interventions. In general, quantile regression provides a robust and comprehensive view of the relationship between covariates and the response variable, leading to many important applications: medicine (Kottas and Gelfand, 2001), finance (Bassett and Chen, 2002), and environmental studies (Pandey and Nguyen, 1999), among many others.
Broadly, there are several important components and considerations in quantile regression. First, quantile-specific linear coefficient estimates are obtained to detect potentially heterogeneous effects, $\beta_{j}(\tau)\neq\beta_{j}(\tau^{\prime})$ for some $\tau\neq\tau^{\prime}$, including both magnitude and direction. Second, quantile-specific uncertainty quantification provides important context for these coefficients and facilitates comparisons across both quantiles and variables. Third, when $p$ is moderate or large, quantile-specific subset selection provides more parsimonious summaries and identifies the most impactful covariates across the distribution $Y\mid\boldsymbol{x}$. Each of these targets must simultaneously respect the fundamental smoothness across quantiles: estimates, uncertainties, and selections should be similar for adjacent quantiles. Finally, the algorithms that deliver these results should be scalable in both the number of observations $n$ and the number of covariates $p$.
We first introduce our motivating application, and then review existing Bayesian and frequentist methods for quantile regression–both to showcase the successes in this area and to highlight the need for methodological advances, especially for our data analysis.
1.1 Social Stressors, Environmental Exposures, and Childhood Educational Outcomes
Childhood educational outcomes are affected by social stressors (e.g., poverty, structural racism, high unemployment) and environmental exposures (e.g., poor air quality, lead exposure), which often cumulate in the same communities (Miranda et al., 2007; Bravo et al., 2024). Thus, there is urgency to quantify and characterize these effects and initiate well-informed policy interventions, especially given the strong links between educational attainment and adult health. However, far less is known about the differential impacts of social stressors and environmental exposures on low-, medium-, and high-achieving students. Such information is critical for understanding the factors that shape educational outcomes and for recommending policy interventions, especially to support at-risk students.
@@ 1.1 Social Stressors, Environmental Exposures, and Childhood Educational Outcomes
There are several important aspects and challenges in this analysis. First, EoG test scores are not merely proxies for educational achievement; more directly, they help determine whether students are promoted to the next grade, participate in programming for intellectually gifted students, or receive access to specialized support and resources. Clearly, it is important to differentiate among the factors that impact low-, medium-, and high-achieving students. Previous analyses of these data have exclusively used (homoscedastic) mean regression for estimation and variable selection (Kowal et al., 2021; Kowal, 2022; Bravo et al., 2022), which is not capable of estimating heterogeneous, quantile-specific effects. Second, the collection of potential variables (Table 1) is moderately large with substantial correlations among the variables (see the supplementary material, Section F.1), since many of the social and environmental exposures appear concurrently. Consequently, subset selection is especially useful, but must be i) quantile-specific, ii) capable of handling correlations among variables, and iii) scalable in $n$ and $p$. Finally, because of the significant policy implications, it is imperative that our numerical summaries are interpretable, robust, and accompanied by full uncertainty quantification. Because our quantile regression analysis will enumerate estimates (effect directions and magnitudes), uncertainties, and selected subsets across many quantiles, this requires stability (or smoothness) and probabilistic coherency across quantiles.
@@
RI Racial residential isolation, at time of EoG test